8 January 2014 – Wednesday

LDAP:

  • Multiple update issues.  Three former students… dance x 3.  Two employees – one a student, one that matches an existing former student but has a different last name.  Sent query to HR about both of the employee issues.  I think I found matches, but I need confirmation.
  • HR responded to yesterday’s query that I had found the right person.  I merged the PeopleSoft information with the existing LDAP entry.
  • ou/title attributes treated as if they do not have matching rules (replace only to preserve value order) rolled onto production server.  Updated 289 entries (out of 8,653) that had pre-existing order problems.

Backups:

  • One of COM/IS’ systems failed to resolve (DNS) last night.  Checking with them whether it has been turned off or something else happened.
    • Received word back that yes, that one is gone and so is one other.  Both removed from backups.
  • Request for three new client definitions.  Asked clarifying questions.  Defined the three new clients, each on a different weekend.
  • Too many tapes are sitting there for more than a month and they should have recycled themselves by now.  Investigation started.
    • Three of the fifty-nine volumes do not have anything that is not recyclable on them.
    • Relabeled those three.  We’ll see what that does to the list.
  • Discovered a typo in the iptables configuration file on stornode2.  This was preventing iptables from running.  Fixed typo and got iptables working again.  Doh!  This is fallout from the crap that was happening with NFS kernel hangs over the Christmas break (I think).  The typo was a case issue in a chain name (iscsi-input vs ISCSI-INPUT).  This was found and fixed while testing the new zoo system backups.
  • Clone status:
    • Tape Clones: Completed
    • COM/IS Clones: Working on the final 2508.6 GB
    • Disk Clones: Working on 912.5 GB, 36,867.5 GB left.

Tags:

Leave a Reply

You must be logged in to post a comment.